(i)Find the integral of f(x,y ) x^4+y^2 over the region bounded by y=x , y= 2x and x=2

(ii)Find the surface area of the portion of the paraboloid z=25–x^2–y^2 which lies

above the xy -plane.

(iii)Locate and classify the stationary points of the function

f( x,y ) x^2+y^2–6xy +6x +3y –4

(iv)Check whether the following functions are homogeneous or not.

(a)x/y +3y/2x +sin√(x/y)

(b)x^4 +4x^2 +y^2)x^2


(v)Evaluate f(xy )at a point (x,y)for the function f defined by

f(x,y)=x^5 +10x^3 y^3 +8y^4

Verify that the function f satisfies the requirements of Schwarz’s theorem and

hence evaluate f(x,y)
1
Expert's answer
2020-03-06T11:13:23-0500

(i)


"\displaystyle\int_{0}^2\displaystyle\int_{x}^{2x}(x^4+y^2)dydx=\displaystyle\int_{0}^2\bigg[x^4y+{y^3 \over 3}\bigg]\begin{matrix}
 2x \\
 x
\end{matrix}dx=""=\displaystyle\int_{0}^2\bigg(x^4(2x)+{(2x)^3 \over 3}-x^4(x)-{x^3 \over 3}\bigg)dx=""=\displaystyle\int_{0}^2\big(x^5+{7x^3 \over 3}\big)dx=\big[{x^6 \over6}+{7x^4 \over 12}\big]\begin{matrix}
 2 \\
 0
\end{matrix}=""={2^6 \over 6}+{7(2)^4 \over 12}-0=20"

"\displaystyle\int_{0}^2\displaystyle\int_{x}^{2x}(x^4+y^2)dydx=20"


(ii) "z=25-x^2-y^2"

The region "E" in the xy-plane is the disk "0\leq x^2+y^2\leq25" (disk or radius 5 centered at the origin).

For this problem, "z=f(x, y)=25-x^2-y^2."

"f_x=-2x, f_y=-2y"

"\sqrt{1+(f_x)^2+(f_y)^2}=\sqrt{1+(-2x)^2+(-2y)^2}=\sqrt{1+4x^2+4y^2}"

The region "E=\{(x,y)|x^2+y^2\leq25\}" is "\{(r,\theta|0\leq r\leq5, 0\leq \theta\leq2\pi\}" in polar coordinates.

"\sqrt{1+4x^2+4y^2}=\sqrt{1+4r^2}"

Hence, the surface area "A" is given by


"A=\int\int_{E}\sqrt{1+(f_x)^2+(f_y)^2}dxdy=""=\displaystyle\int_{0}^{2\pi}\displaystyle\int_{0}^{5}\sqrt{1+4r^2}rdrd\theta=""=\displaystyle\int_{0}^{2\pi}\bigg[{(1+4r^2)^{3/2} \over 12}\bigg]\begin{matrix}
 5 \\
 0
\end{matrix}d\theta=""={2\pi \over 12}\big((1+4(5)^2)^{3/2}-1\big)={\pi \over 6}\big((101)^{3/2}-1\big)"

(iii)

"f(x,y)=x^2+y^2-6xy+6x+3y-4"

Partial derivatives

"f_x=2x-6y+6"

"f_y=2y-6x+3"

To find the critical points, we solve

"f_x=0=>2x-6y+6=0"

"f_y=0=>2y-6x+3=0"


"x=-{111 \over 16}, y=-{21 \over 16}"

We find the critical point "(-{111 \over 16},-{21 \over 16})"

To find the nature of the critical point, we apply the second derivative test. We have

"A=f_{xx}=2, B=f_{xy}=-6, C=f_{yy}=2"

At the point "(-{111 \over 16},-{21 \over 16})," we have


"D=AC-B^2 =2(2)-(-6)^2=-32<0"

Then the point "(-{111 \over 16},-{21 \over 16})" is a saddle point.

(iv)Check whether the following functions are homogeneous or not.

(a)


"f(x, y)={x \over y}+{3y \over 2x}+\sin{\sqrt{{x \over y}}}""f(tx, ty)={tx \over ty}+{3ty \over 2tx}+\sin{\sqrt{{tx \over ty}}}=""={x \over y}+{3y \over 2x}+\sin{\sqrt{{x \over y}}}=t^0 f(x, y)"

A function "f(x, y)" is homogeneous.


(b)


"f(x, y)=x^4+4x^2+{y^2 \over x^2}""f(tx, ty)=(tx)^4+4(tx)^2+{(ty)^2 \over (tx)^2}=t^4x^4+4t^2x^2+{y^2 \over x^2}"


A function "f(x,y)" is not homogeneous.


(v)


"f(x,y)=x^5+10x^3y^3+8y^4"

Evaluate "f_{xy}" at a point (x,y)


"f_x(x,y)=5x^4+30x^2y^3""f_{xy}(x,y)=90x^2y^2"

Verify that the function "f" satisfies the requirements of Schwarz’s theorem and hence evaluate f(x,y)

Since "90x^2y^2" is a polynomial, "f_{xy}(x,y)" is a continuous function. 

Further, "f_y(x,y)=30x^3y^2+32y^3" exists. Hence "f" satisfies the conditions of Schwarz's theorem and so 


"f_{yx}(x,y)=f_{xy}(x,y)=90x^2y^2"